# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

### v2

@@ III-A4 Dataset Summary and Statistics
The final dataset contains a total of 3,347 PennyLane code samples, as shown in Table IV.
TABLE IV: Dataset Composition.
Source Samples Collected
Unofficial GitHub repositories 1,952
GitHub (PennyLaneAI repository + others) 1,321
Official PennyLane Documentation 53
Quantum Computing Books 21
Total Dataset Size 3,347
Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for fine-tuning large language models (LLMs) for PennyLane-based quantum programming.
@@ III-A5 Ensuring Data Integrity and Diversity
To maintain dataset integrity, we applied multiple quality control measures:
@@ III-A6 Implications for AI-Driven PennyLane Code Generation
  • Improving Model Performance: High-quality annotations ensure that models understand both syntax and context.
  • Facilitating Explainable AI: Models trained on our dataset can generate explanatory comments alongside code.
@@
TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework.
Component Function Description
DirectoryLoader Document Loader Loads PennyLane dataset samples as text files from a directory.
Chroma Vector Database Stores embeddings for retrieval using similarity search.
OpenAIEmbeddings Embedding Model Converts text into vector embeddings for indexing and retrieval.
ChatPromptTemplate Prompt Formatting Defines structured prompts for AI models to ensure consistency.
create_retrieval_chain Retrieval Pipeline Creates a pipeline that integrates stored embeddings for improved responses.
ChatOpenAI OpenAI LLM Interface Connects to OpenAI models (GPT-4o-mini) for response generation.
ChatAnthropic Anthropic LLM Interface Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses.
TABLE VI: LLMs Used in the RAG-Based Evaluation Framework.
Model Parameters Architecture Provider Usage in Pipeline
Qwen2.5-7B-Instruct-Turbo 7B Causal Decoder-Only Transformer Alibaba Cloud Generates PennyLane code
GPT-4o Mini Not disclosed Causal Decoder-Only Transformer OpenAI Evaluates and scores generated responses
Claude 3.5 Sonnet Not disclosed Causal Decoder-Only Transformer Anthropic Alternative model for response generation
@@ III-B1 Handling Variable-Length Code Samples
The collected PennyLane code samples exhibit varying lengths after tokenization. Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$. Following standard practice, left padding was applied to align shorter samples to $N$ tokens, ensuring batch processing efficiency and minimizing computational overhead.
@@ III-B2 Padding Strategies and Tokenizer Variations
  • Left vs. Right Padding: Left padding is commonly used in autoregressive models [52], ensuring that padding tokens do not interfere with causal attention mechanisms.
In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, and Claude 3.5 Sonnet. This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining autoregressive consistency.
@@ III-B3 Applying Attention Masking
To improve transformer efficiency, an attention mask [53] was applied post-padding. This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data. Table III represents the attention mask matrix applied to a batch of tokenized samples. Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
@@ III-C1 Retrieval-Augmented Generation (RAG) Framework
To assess the impact of context-aware prompting on LLM-based quantum code generation, we adopt a Retrieval-Augmented Generation (RAG) framework. This method enhances generation quality by incorporating relevant external context retrieved from our curated PennyLang dataset. RAG helps address common shortcomings in LLMs, such as hallucinations and outdated or incorrect syntax, by anchoring their output in concrete examples from real-world PennyLane programs. This results in improved syntactic validity, functional correctness, and alignment with PennyLane best practices.
III-C2 RAG Implementation using LangChain
Our RAG pipeline is implemented using the LangChain framework, which provides a modular infrastructure for document loading, vector embedding, retrieval, and prompt formatting. The dataset—formatted in an instruction-response structure—is first embedded using the OpenAI Embeddings module. The resulting vectors are stored in a Chroma database for efficient retrieval.
@@ III-C2 RAG Implementation using LangChain
When a user query is submitted (e.g., a quantum programming task or challenge description), we apply Maximum Marginal Relevance (MMR) search to extract the top relevant examples from the dataset. These examples are concatenated into a structured prompt and fed into the LLM for code generation. To ensure fairness and evaluate RAG’s effectiveness, we generate responses both with and without retrieved context using the same query.
III-C3 RAG Evaluation
To benchmark the performance of RAG-augmented LLMs, we evaluate three state-of-the-art models: GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5-7B-Instruct-Turbo. Each model is tested on 29 quantum code generation tasks sampled from the QHack challenge dataset. For each query, both a baseline (non-RAG) and a RAG-enhanced response is generated. We then use GPT-4o to evaluate the outputs according to a rubric comprising three criteria:
  • Functionality: Does the code produce the correct quantum behavior?
  • Syntax: Is the code valid and executable in PennyLane?
  • Modularity: Is the solution well-organized and reusable?
Figure 4: Performance of GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5 with and without RAG. Mean scores (±1 SD) over 29 trials are shown for each model. Blue: RAG; Orange: Non-RAG.
Figure 4 summarizes the outcome of these evaluations. RAG-enhanced prompting consistently improves model performance across all criteria. On average, GPT-4o Mini improves by 11.67%, Claude 3.5 by 7.69%, and Qwen 2.5 by 14.38%. These gains highlight how retrieval provides critical context, especially for complex or multi-step quantum circuits.
III-C4 GraphRAG for Advanced Retrieval
To further improve retrieval precision beyond vector-based similarity, we extend our RAG framework using GraphRAG. While the baseline RAG implementation relies solely on Maximum Marginal Relevance (MMR) vector retrieval, GraphRAG incorporates structural relationships between code samples to guide context selection. In this method, the PennyLang dataset is organized into a graph where nodes represent code samples and edges represent functional or structural similarity—such as shared quantum gates, observables, and template usage patterns.
When a query is received, GraphRAG identifies a subgraph around the most semantically similar samples and performs hybrid retrieval by combining graph neighborhood traversal with traditional vector similarity. This enables the selection of more functionally aligned and topologically relevant code snippets, resulting in higher quality context for LLMs.
We evaluate GraphRAG using the same task set as our baseline experiments but restrict the evaluation to the GPT-4o Mini model due to its superior baseline performance. The model was tested on 146 PennyLang test tasks in both no-RAG and GraphRAG settings. The correctness of the generated outputs was determined by whether the solution passed the associated unit tests.
TABLE VII: Performance of GPT-4o Mini: No-RAG vs. GraphRAG (146 tasks).
Metric No-RAG GraphRAG
Correct Solutions 30 85
Incorrect Solutions 116 61
Total Samples 146 146
Accuracy (%) 20.55% 58.22%
As shown in Table VII, the GraphRAG approach yields a dramatic improvement in accuracy—from 20.55% without retrieval to 58.22% with graph-enhanced retrieval. This threefold increase highlights the value of graph-based structural reasoning in quantum programming tasks, particularly for problems involving complex circuit construction or uncommon PennyLane APIs.
Together, these results establish GraphRAG as a significant enhancement to conventional RAG pipelines and demonstrate its practical utility in improving the performance of LLMs on challenging real-world quantum programming benchmarks.
@@ IV-B Instruction and Response Length Distributions
Figure 3 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset. Instruction lengths, as shown in the left plot, follow a normal distribution with a peak around 35 words. The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare. In contrast, response lengths exhibit a right-skewed distribution, as shown in the middle plot. Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs. The instruction-response correlation, illustrated in the rightmost scatter plot, reveals a positive correlation between instruction and response length. While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
@@ IV-C Feature Group Usage Across Quantum Domains
Figure 6: Feature Group Usage Across Domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories, with color intensity reflecting count.
@@ IV-E Insights and Implications
The findings from Figures 3, 5, 7, and 6 offer several key insights:
@@
V Case Studies
The case studies compare base LLM-generated PennyLane code with RAG-enhanced implementations, showcasing notable improvements in functionality, syntax, and modularity across different quantum programming tasks.
In the first task, implementing a 1-qubit parameterized RX gate, the base model introduced redundant imports, suboptimal measurement choices, and missing hardware-specific optimizations. The RAG-optimized implementation addressed these issues by incorporating hardware-aware device initialization, optimized measurement practices, and type annotations, achieving a perfect 10/10 functionality score.
The second task, developing an RX gate circuit with optional parameter binding, revealed static parameter handling and the absence of deferred execution in the base LLM output. The RAG-enhanced version introduced flexible parameter binding, probabilistic measurement strategies, and deferred execution patterns, improving adaptability and modularity.
The third case study, constructing a variational quantum classifier using StronglyEntanglingLayers, further demonstrated the superiority of RAG-based generation. The base LLM solution lacked adaptive training integration, efficient optimization, and TensorFlow-Keras support, requiring manual gradient updates. In contrast, the RAG-optimized model seamlessly integrated TensorFlow, Keras, and the Adam optimizer, significantly enhancing training efficiency and parameter optimization.
Overall, across all tasks, RAG-assisted code generation consistently outperformed the base LLM, producing more structured, modular, and functionally superior PennyLane code. This highlights the effectiveness of retrieval-augmented approaches in enhancing quantum programming workflows.
Task: Implement 1-qubit parameterized RX gate using PennyLane
Base LLM Implementation (Without RAG) Limitations Identified: Unnecessary numpy import Suboptimal measurement choice (expectation value vs state) Missing hardware-specific optimizations 1import pennylane as qml 2from pennylane import numpy as np # Redundant import 3 4dev = qml.device(’default.qubit’, wires=1) 5 6@qml.qnode(dev) 7def circuit(theta): 8 qml.RX(theta, wires=0) 9 return qml.expval(qml.PauliZ(0)) # Non-optimal measurement Listing 1: Non-RAG Implementation Evaluation Results: Functionality: 9/10 (Valid but limited measurement) Syntax: 9/10 (Redundant imports) Modularity: 8/10 (Missing type hints)
RAG-Enhanced Implementation Retrieved Context: Hardware-aware device initialization (sample_1137.txt) State measurement best practices (sample_214.txt) Testing methodologies (sample_23.txt) 1import pennylane as qml 2 3dev = qml.device("lightning.qubit", wires=1) # Hardware-optimized 4 5@qml.qnode(dev) 6def circuit(theta: float): # Type hinting 7 qml.RX(theta, wires=0) 8 return qml.state() # Full state measurement Listing 2: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Complete state analysis) Syntax: 10/10 (Clean imports) Modularity: 9/10 (Type annotations)
Task: Create 1-qubit RX gate circuit with optional parameter binding
Base LLM Implementation (Without RAG) Limitations Identified: Static parameter handling Missing deferred parameterization Limited measurement flexibility 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 if theta: # Binary parameter check 6 qml.RX(theta, wires=0) 7 return qml.expval(qml.PauliZ(0)) # Fixed measurement Listing 3: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (No variable binding) Syntax: 9/10 (Valid but rigid) Modularity: 8/10 (No late binding support)
RAG-Enhanced Implementation Retrieved Context: Variable parameter techniques (sample_1137.txt) Probabilistic measurement strategies (sample_214.txt) Deferred execution patterns (sample_23.txt) 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 param = theta if theta else qml.variable(’theta’) 6 qml.RX(param, wires=0) 7 return qml.probs(wires=0) # Full probability measurement Listing 4: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Flexible parameter binding) Syntax: 9/10 (Clear variable handling) Modularity: 9/10 (Supports late binding)
Task: Implement a Variational Quantum Classifier with Strongly Entangling Layers
Base LLM Implementation (Without RAG) Limitations Identified: Uses a static quantum circuit with fixed parameters, lacking adaptive training integration. Lacks integration with TensorFlow/Keras, requiring a manual optimization loop. Uses Gradient Descent Optimizer instead of the Adam optimizer, which may slow convergence. 1import pennylane as qml 2import numpy as np 3from sklearn.datasets import make_classification 4from sklearn.model_selection import train_test_split 5from sklearn.preprocessing import StandardScaler 6from sklearn.metrics import accuracy_score 7 8# Generate a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Scale the features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define the quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Define the quantum circuit 21@qml.qnode(dev) 22def circuit(params, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.StronglyEntanglingLayers(params, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 5: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (Works but lacks adaptive training integration) Syntax: 9/10 (Code is valid but manual optimization is not ideal) Modularity: 8/10 (Training and inference should be better structured)
RAG-Enhanced Implementation Retrieved Context: TensorFlow-Keras integration for efficient optimization. Adam optimizer instead of manual updates for improved convergence. Improved quantum circuit design based on best practices. 1import pennylane as qml 2import numpy as np 3import tensorflow as tf 4from sklearn.datasets import make_classification 5from sklearn.model_selection import train_test_split 6from sklearn.preprocessing import StandardScaler 7 8# Create a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Standardize features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Quantum circuit with Strongly Entangling Layers 21@qml.qnode(dev) 22def quantum_circuit(weights, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.templates.StronglyEntanglingLayers(weights, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 6: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Seamless Keras integration for training) Syntax: 10/10 (Concise and efficient TensorFlow pipeline) Modularity: 9/10 (Well-structured training process)

### v3

@@ III-A4 Dataset Summary and Statistics
The final dataset containing 3,347 PennyLane code samples is shown in Table III.
TABLE III: Dataset Composition
Source Samples Collected Unofficial GitHub repositories 1,952 GitHub (PennyLaneAI repository + others) 1,321 Official PennyLane Documentation 53 Quantum Computing Books 21 Total Dataset Size 3,347
Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for enhancing LLMs for PennyLane-based quantum programming.
@@ III-A5 Ensuring Data Integrity and Diversity
To keep dataset integrity, we applied multiple quality control measures:
@@ III-A6 Implications for AI-Driven PennyLane Code Generation
  • Improving Model Performance: High-quality annotations ensure models understand both syntax and context.
  • Facilitating Explainable AI: Models trained on PennyLang can generate explanatory comments alongside code.
@@ III-B1 Handling Variable-Length Code Samples
The collected PennyLane code exhibit varying lengths after tokenization. Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$. Following standard practice, left padding was applied to align shorter samples to $N$ tokens, ensuring batch processing efficiency and minimizing computational overhead.
@@ III-B2 Padding Strategies and Tokenizer Variations
  • Left vs. Right Padding: Left padding is commonly used in autoregressive models [50], ensuring that padding tokens do not interfere with causal attention.
In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, Claude 3.5 Sonnet, and LLaMa 4. This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining auto-regressive consistency.
@@ III-B3 Applying Attention Masking
To improve transformer efficiency, an attention mask [51] was applied post-padding. This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data. Table IV shows the attention mask matrix applied to a batch of tokenized samples. Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
TABLE IV: Tokenized Input and Attention Masking.
Token Mask Token Mask Token Mask "<pad>" 0 "def" 1 "circuit" 1 "<pad>" 0 ":" 1 "(" 1 ")" 1 "return" 1 "qml.expval" 1
@@ III-C1 Retrieval-Augmented Generation (RAG) Framework
To improve PennyLane code generation by reducing hallucinations and outdated syntax—and to better align outputs with PennyLane best practices—we adopt a retrieval-augmented generation (RAG) [52] approach. We implement the pipeline in LangChain with modular components for document ingestion, vector embedding, retrieval, and prompt construction. The PennyLang dataset, formatted as instruction–response pairs, is embedded using OpenAIEmbeddings and stored in a Chroma vector database for efficient similarity search. At inference, user queries are embedded, relevant context is retrieved, and the combined prompt is sent to the target LLM. Table V summarizes the key PennyLang RAG framework components, including their role in dataset loading, embedding storage, prompt formatting, retrieval chaining, and interfacing with multiple LLM backends.
TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework
PennyLang RAG Framework Components DirectoryLoader (Document Loader)    $\triangleright$ Loads PennyLane dataset samples as text files from directory Chroma (Vector Database)    $\triangleright$ Stores embeddings for retrieval using similarity search OpenAIEmbeddings (Embedding Model)    $\triangleright$ Converts text into vector embeddings for indexing and retrieval ChatPromptTemplate (Prompt Formatting)    $\triangleright$ Defines structured prompts for AI models to ensure consistency create_retrieval_chain (Retrieval Pipeline)    $\triangleright$ Creates pipeline integrating stored embeddings for improved responses ChatOpenAI (OpenAI LLM Interface)    $\triangleright$ Connects to OpenAI models (GPT-4o-mini) for response generation ChatAnthropic (Anthropic LLM Interface)    $\triangleright$ Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses
@@
When a user query is submitted (e.g., a quantum programming task or challenge description), we use Maximum Marginal Relevance (MMR) search to retrieve the most relevant examples from the dataset. The retrieved examples are concatenated into a structured prompt and passed to the LLM for code generation. For a fair assessment of RAG’s impact, we generate outputs for each query both with and without retrieved context under identical conditions.
@@ IV-B Instruction and Response Length Distributions
Figure 4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset. Instruction lengths, as shown in left plot, follow a normal distribution with a peak around 35 words. The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare. In contrast, response lengths exhibit a right-skewed distribution, as shown in the middle plot. Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs. The instruction-response correlation, illustrated in the rightmost plot, reveals a positive correlation between instruction and response length. While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
@@ IV-C Feature Group Usage Across Quantum Domains
Figure 5: Feature group usage across domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories (QML, Quantum Hardware, Quantum Chemistry and Optimization), with color intensity reflecting count.
TABLE VI: Evaluation of Qwen 2.5 7B, LLaMa 4 Maverick 17B, GPT-4o-mini, and Claude 3.5 Sonnet on PennyLane
Metric Qwen 7B LLaMa 4 GPT-4o-mini Claude 3.5 Sonnet Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Total Notebooks 1320 1320 1320 1320 Working Notebooks 467 495 480 95 901 875 895 799 712 780 755 625 899 911 856 1001 Success Rate (%) 41.66 45.07 43.18 8.71 84.84 84.84 82.17 78.78 80.30 84.84 81.81 85.60 89.01 92.44 87.50 95.07 Pass@1 87 100 88 19 179 171 179 150 144 158 149 134 187 182 170 194 Pass@2 99 109 103 21 198 202 199 178 171 189 175 175 210 218 203 230 Pass@3 104 113 109 21 212 214 209 191 190 207 195 199 223 232 214 238 Pass@4 109 117 110 22 216 217 212 202 202 219 209 215 230 242 230 246 Pass@5 110 119 114 23 224 224 217 208 212 224 216 226 235 244 231 251
@@ IV-E Insights and Implications
Findings from Figures 4, 3, 6, and 5 offer several insights:
@@
V Evaluation
To assess the effectiveness of our proposed dataset for quantum code generation, we designed a benchmark suite comprising 264 test cases, each tailored to evaluate various PennyLane functionalities and quantum programming concepts. Each test was used as a prompt within an RAG pipeline and evaluated across both open-source and commercial LLMs. The evaluation utilized the Pass@5 metric, where each model generates five candidate solutions per test, and success rate (%) is computed as $\text{pass@5}/{264}\times 100$, where pass@5 indicates the number of test cases for which at least one correct solution was generated among the five attempts. In total, 1320 notebooks were generated and executed per model (264 tests × 5 completions), ensuring a rigorous, code-level assessment of correctness. Table VII lists the details for evaluated models.
TABLE VII: LLMs Used in the RAG-Based Evaluation Framework
Large Language Models in RAG Pipeline Qwen2.5-7B-Instruct-Turbo (Alibaba Cloud)    $\bullet$ Parameters: 7B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Generates PennyLane code LLaMa 4 Maverick (Meta)    $\bullet$ Parameters: 17B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Code generation (planned evaluation) GPT-4o Mini (OpenAI)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Evaluates and scores generated responses Claude 3.5 Sonnet (Anthropic)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Alternative model for response generation
Open-Source Models.
Given that the dataset was primarily developed to support open-source development, our core evaluation focused on Qwen 7B and LLaMa 4. These models showed clear improvements when used with our dataset. Notably, RAG with 75% context coverage often outperformed the full-context setting. This suggests that a moderate amount of relevant context may reduce noise and improve grounding, while excessive retrieval might introduce redundancy or irrelevant tokens.
Commercial Models.
For comparison, we also tested GPT-4o-mini and Claude 3.5. As expected, these commercial models demonstrated high performance even without RAG. However, the added context from our dataset provided limited benefit. This may be attributed to their broad pretraining, which likely includes similar material (e.g., PennyLane documentation or equivalent quantum code examples), reducing the marginal gain from external retrieval.
Discussion.
The evaluation reveals important interactions between dataset design, model capacity, and retrieval strategies. Open-source models such as Qwen 7B and LLaMa 4 showed notable improvements when augmented with PennyLang, confirming that domain-specific RAG can effectively compensate for limited pretraining in quantum programming. Interestingly, the 75% context setting consistently outperformed the full-context retrieval, indicating that focused and moderately sized context windows improve model grounding by avoiding redundancy and context saturation.
In contrast, commercial models like GPT-4o Mini and Claude 3.5 achieved high performance across all settings, with only marginal benefits from added retrieval. This behavior is likely due to their extensive pretraining on diverse codebases, which may already include material similar to what our dataset provides. Consequently, the impact of external retrieval is diminished and can even introduce noise if not precisely targeted. These findings emphasize the importance of retrieval quality over quantity in RAG-based code generation. The consistent advantage of partial over full context suggests that carefully curated, relevant examples are more beneficial than exhaustive inclusion.
Finally, by evaluating 1,320 generated notebooks per model, we ensure that our conclusions are grounded in executable and functionally correct outputs, a crucial criterion in the domain of quantum programming. These results underscore the utility of PennyLang in enhancing open-source LLM performance while providing insights into optimal RAG configurations for technical code generation tasks.